Giving a test to a group of students, the grades and gender are summarized below
A B C Total
Male 8 2 5 15
Female 19 14 9 42
Total 27 16 14 57
If one student was chosen at random, determine the following probabilities. Write your answers as reduced fractions.
P
(Student was male) =

P
(Student was female) =

P
(Student was male and got an "A") =

P
(Student was female and got a "B") =

P
(Student got a "C") =


Answer by math_tutor2020(3817)


Table:
ABCTotal
Male82515
Female1914942
Total27161457


Use the table to answer the following shown below.
The final answers are marked in red.

P(male) = (number of males)/(number total)
P(male) = 15/57
P(male) = (3*5)/(3*19)
P(male) = 5/19

P(female) = (number of females)/(number total)
P(female) = 42/57
P(female) = (3*14)/(3*19)
P(female) = 14/19
Or you could say
P(male) + P(female) = 1
P(female) = 1 - P(male)
P(female) = 1 - 5/19
P(female) = 19/19 - 5/19
P(female) = (19 - 5)/19
P(female) = 14/19

P(male & got an A) = (number of males who got an A)/(number total)
P(male & got an A) = 8/57

P(female & got an B) = (number of females who got a B)/(number total)
P(female & got an B) = 14/57

P(student got a C) = (number who got a C)/(number total)
P(student got a C) = 14/57